### v3

Abstract
In recent decades, several assistive technologies have been developed to improve the ability of blind and visually impaired individuals to navigate independently and safely. At the same time, simultaneous localization and mapping (SLAM) techniques have become sufficiently robust and efficient to be adopted in developing these assistive technologies. We present the first systematic literature review of 54 recent studies on SLAM-based solutions for blind and visually impaired people, focusing on literature published from 2017 onward. This review explores various localization and mapping techniques employed in this context. We discuss the advantages and limitations of these techniques for blind and visually impaired navigation. Moreover, we examine the major challenges described across studies. We explain how SLAM technology offers the potential to improve the ability of visually impaired individuals to navigate effectively. Finally, we present future opportunities and challenges in this domain.
Index Terms:
Navigation, SLAM, systematic literature review, visually impaired
[…]
=-21pt
I Introduction
In recent decades, there has been increasing research interest in developing assistive technologies to enhance spatial navigation for blind and visually impaired (BVI) individuals. In most cases, the main goal is to guide and assist BVI people in navigating safely in unknown environments without the help of a sighted assistant. Navigation is a complex task; it requires finding an optimal path to the desired destination, perceiving the surroundings, and avoiding obstacles. Crucially, all of these functionalities need to accurately localize the BVI user in the environment. There are several approaches for localization, such as the global positioning system (GPS), radio frequency identification (RFID), and simultaneous localization and mapping (SLAM) [1, 2]. Each has advantages and challenges and is used in different applications.
GPS is a localization technique employed in outdoor scenarios owing to its affordability to the end user, wide coverage of the Earth, and ease of integration with other technologies. However, this technique suffers from limitations like satellite signal blockage, inaccuracy, and signal loss caused by weather conditions, walls, and other obstacles [2]. Approaches based on RFID utilize small, low-cost tags for localization. To localize an agent, a set of RFID tags must be installed in the environment [1]. Although localization can be accurately performed using an RFID scheme, taking advantage of this technique requires a pre-installed infrastructure.
[…]
SLAM uses diverse types of sensors to determine an agent’s position, location, and velocity and detect and avoid obstacles, even in a dynamically changing unknown environment. This technique uses infrared (IR) sensors, acoustic sensors, RGB cameras, inertial measurement units (IMUs), ultrawide-band (UWB), LiDAR, RADAR, and RGB-D sensors [4].
The collaborative effort between the front- and back-ends empowers SLAM to provide a robust and real-time spatial understanding, making it a valuable tool for various applications. The SLAM community has made tremendous strides over the past 35 years, developing large-scale practical applications and seeing a steady transition of this technology into the industry [3]. Cadena et al. [3] divided the life cycle of SLAM into three periods: classical age (1986-2004), algorithmic analysis age (2004–2015), and robust-perception age (2015-present). The evolution of portable computation and the availability of low-cost, highly accurate, and lightweight sensors such as cameras and IMUs make them appropriate for pedestrian navigation. By exploiting these advances, many researchers have recently adopted SLAM to develop assistive technology demonstrators to help BVI people navigate unknown environments.
Since the first electronic travel aids (ETAs) emerged approximately 70 years ago, the development of navigation devices to guide BVI people through indoor and/or outdoor environments has remained a challenge and a key concern for researchers [5]. From traditional to deep-learning-based navigation approaches, researchers have always faced challenges ranging from technical issues to the limitations of user capabilities. As BVI navigation approaches must improve real-time performance while reducing the size, weight, energy cost, and overall price of the assistive system, these studies have put a lot of effort into coping with constraints in computational issues, sensory equipment, and portable devices. They also need to provide solutions to calculate the precise position and orientation of the user in a real-time manner. However, the challenges of different scenarios, including complex and cluttered environments, noisy environments, and large spaces, must be considered.
Furthermore, efficient and reliable obstacle detection in both indoor and outdoor environments has always been a concern. In this regard, other challenges include identifying static and dynamic obstacles, predicting the risk of collision, understanding moving objects’ motion and estimating their speed, detecting small objects, and identifying obstacles at different levels of the user’s body, from drops in terrain to head level. In addition, an intuitive, user-friendly, low-cognitive-load method to provide accurate and sufficient environmental information to the user is also considered an important research target. These methods should be improved to provide adjustable and customized feedback on demand for different users.
Moreover, assistive technology should provide user safety and independence, hands-free operations, decreased effort, and backup in the case of system failure. In addition to the aforementioned challenges, deep-learning-based solutions also have special issues, such as designing lightweight neural network architectures to reduce computational expense and provide sufficient data for the training and validation of the models.
I-A Related Work
Thus far, many reviews have been conducted on assistive technologies developed for BVI navigation. Several studies reviewed walking assistance systems [14, 7, 8, 9, 10, 11, 12, 5, 13] provided a detailed classification of the developed approaches. [12] categorized walking assistants into three groups: sensor-based, computer vision-based, and smartphone-based. The authors explained the technologies used and inspected each approach, and evaluated some important parameters of each approach, such as the type of capturing device, type of feedback, working area, cost, and weight. The work by [14] introduced techniques and technologies designed to assist visually impaired individuals in their mobility and daily lives. This comprehensive review analyzes multiple mobility-assistive technologies that are suitable for indoor and outdoor environments. It offers insights into the various feedback methods employed by assistive tools based on recent technologies. In addition, [7] reviewed wayfinding devices used by visually impaired individuals in real-world scenarios. This review aimed to provide a comprehensive exploration of the various aids employed for navigation while assessing their perceived efficacy.
Some studies focused on indoor navigation for BVI users [15, 16, 17, 18, 19, 20, 21, 22] and some focused on computer vision-based navigation systems [15, 23, 24, 25, 26, 27]. Among these studies, [15] conducted a systematic literature review of state-of-the-art computer vision-based methods used for indoor navigation. The authors described the advantages and limitations of each solution under review, and included a brief description of each method. Furthermore, [21] comprehensively examined existing methods and systems developed within the domain of assistive technology, with a specific focus on addressing the unique needs and challenges faced by the visually impaired. This study places strong emphasis on evaluating methods that have practical applications in enhancing the lives of visually impaired individuals.
Several review papers on wearable navigation systems have also been published [28, 29, 30, 31, 32, 33]. [28] have conducted a systematic review with the primary objectives of analyzing wearable obstacle avoidance electronic travel aids. Their work delves into the strengths and weaknesses of existing ETAs, providing a thorough evaluation of hardware functionality, cost-effectiveness, and the overall user experience.  [29] provided a comprehensive understanding of wearable travel aids by focusing on their designs and usability. Their objectives included surveying the current landscape of travel aid design, investigating key design issues, and identifying limitations and future research directions. Furthermore, [30] conducted a systematic review of the literature on wearable technologies designed to enhance the orientation and mobility of the visually impaired. This review provides valuable insights into the technological characteristics of wearables, identifies feedback interfaces, emphasizes the importance of involving visually impaired individuals in prototype evaluations, and highlights the critical need for safety evaluations. A review by [31] provides a comprehensive review of computer vision and machine-learning-based assistive methods. Existing ETAs are divided into two groups: active systems providing subject localization and object identification, and passive systems providing information about the users’ surroundings using a stereo camera, monocular camera, or RGB-D camera.
Focusing on guide robots, [34] reviewed the multifaceted objectives. Their work included a comparative analysis of the existing robotic mobility aids and state-of-the-art technologies. This review highlights the potential of guide robots to enhance the mobility and independence of the visually impaired.
[35] and [36] reviewed studies with the focus on object detection and recognition. [36] performed a review on object recognition tailored to the needs of visually impaired individuals. This review examines state-of-the-art object detection and recognition techniques, focuses on standard datasets, and emphasizes on the latest advancements. [35] reviewed studies specific for staircase detection systems, primarily designed to facilitate the navigation of visually impaired individuals. The goal of this review is to provide a comprehensive comparative analysis of these systems considering their suitability and effectiveness.
Other similar studies include a survey of inertial measurement units (IMUs) in assistive technologies for visually impaired people [37], a review of urban navigation for BVI people [38], a survey paper that reviewed assistive tools based on white canes [39], and review papers exploring smartphone-based navigation devices [40, 41, 42].  [40] reviewed the multifaceted objectives in the domain of smartphone-based navigation devices. They aimed to provide a comprehensive overview of smartphone use among people with vision impairment, identify research gaps for future exploration, and delve into the use of smartphones by individuals with vision impairment and the accessibility challenges they encounter. To the best of our knowledge, there is no survey paper on SLAM-based navigation systems for BVI people. Our study aims to bridge this gap in literature.
I-B Contribution
This paper presents a systematic literature review (SLR) addressing fundamental questions regarding SLAM-based approaches for BVI navigation. This review provides insights into technological diversity, advantages, limitations, and the potential to address real-world challenges. While recognizing the broad range of potential research questions, such as sensors employed in studies, evaluation techniques, and human–computer interaction for feedback, we narrowed our focus to the four questions outlined in Table II. The primary contributions of this study are as follows:
  • Identification of SLAM Approaches: We systematically identify and categorize the diverse SLAM approaches adopted in the development of assistive systems tailored for visually impaired navigation.
  • Advantages and limitations synthesis: Our study synthesizes the advantages and limitations of these SLAM techniques when applied to BVI navigation.
  • Classification of challenges: We identify and categorize studies that address challenging conditions relevant to SLAM-based navigation systems for the visually impaired.
  • Exploration of the potential for enhancing BVI navigation: We show that SLAM-based solutions have the potential to enhance the mobility and navigation capabilities of individuals with visual impairment.
I-C Paper structure
The remainder of this paper is organized as follows. In Section II, we explain the protocol, methodology, tools, and techniques used to conduct SLR. The findings of our SLR and answers to the SLR research questions are summarized in Section III. Section IV presents the future opportunities and potential advancements in this domain. Finally, Section V concludes the paper.
II SLR methodology
A systematic literature review is one of the most common types of literature review used to collect, review, appraise, and report research studies on a specific topic, adhering to predefined rules for conducting the review [43]. Compared with traditional literature reviews, it provides a wider and more precise understanding of the topic under review [44]. Various guidelines exist for conducting SLR in different research fields such as software engineering [45, 46, 47], computer science [48], information systems [49], planning education and research [50], and health sciences [51, 44]. To conduct this review, we followed the guidelines for conducting systematic reviews proposed by [52]. Figure 1 illustrates our SLR process.
Figure 1: The process of SLR
SLR consists of three key phases: planning, conducting, and reporting the review. We defined our research questions and motivation, keywords, and search string, as well as selection criteria in the planning phase of the SLR. In the conducting phase, we executed searches on digital sources using predefined search strings that were established during the planning stage. We evaluated the quality of the selected papers and extracted relevant data aligned with SLR research questions.
We used the PICOC criteria to identify the key elements that needed to be considered and frame our research questions. PICOC represents Population, Intervention, Comparison, Outcome, and Context [53]. Table I lists the PICOC elements, relevant values, and descriptions of these elements in our study.
TABLE I: Elements of PICOC
Element
Description [53]
Value
Population
The problem or situation
Visually impaired navigation
Intervention
The technology, tool or method under study
SLAM
Comparison (optional)
The technology, tool or method with which the intervention is compared
Outcome
Results that Intervention could produce
Lightweight, affordable, accurate, efficient assistive technology
Context
The specific context of the study
Autonomous mobility of visually impaired people
In this section, we first introduce the tool we used to manage our SLR process and then detail our methodology for conducting our systematic literature review in planning and conducting the review subsections.
II-A SLR Tool
Various tools have been used to conduct systematic literature reviews. Some of them are commercial such as Covidence [^1] https://www.covidence.org/home, DistillerSR [^2] https://www.evidencepartners.com/products/distillersr-systematic-review-software/, and EPPI-Reviewer [^3] https://eppi.ioe.ac.uk/CMS/Default.aspx; and some are free such as Cadima [^4] https://www.cadima.info/, Rayyan555 https://rayyan.qcri.org/welcome, RevMan [^6] https://training.cochrane.org/online-learning/core-software-cochrane-reviews/revman, and Parsifal [^7] https://parsif.al. We used the Parsifal platform to manage the SLR phases. It is an online tool developed to support the process of performing SLR. Parsifal provides researchers with an interface to invite co-authors to collaborate in a shared workspace on the SLR. During the planning phase, this tool assists the authors by addressing the objectives, PICOC, research questions, search strings, keywords and synonyms, selection of sources, and inclusion and exclusion criteria. Parsifal offers tools for creating a quality assessment checklist and data extraction forms. In the conducting phase, this tool helps the authors import the bibtex files and select studies. It assists in identifying and eliminating duplicates among various sources, performing quality assessments, and facilitating data extraction from papers. Finally, it provides a method to document the entire SLR process.
II-B Planning the review
The first step in conducting SLR is to establish a protocol. The protocol outlines the review procedures and ensures replicability. Within the protocol, we formulated our research questions, designed a search strategy, and defined the specific criteria for selecting relevant studies. In addition, we defined a set of criteria presented in Table VI to evaluate the quality of the selected literature. Furthermore, to facilitate the extraction of data in alignment with our research questions, we designed a data-extraction form.
II-B1 Research Questions and Motivation
The SLAM technique is widely used for the navigation of robots, autonomous drones, and self-driving cars, owing to its performance, reliability, and efficiency. Therefore, we reviewed the literature on visually impaired navigation, which was designed based on the SLAM technique. Our aim was to determine the advantages and limitations of employing this technique for visually impaired navigation as well as to identify opportunities for future research. Furthermore, we aimed to explore how extensively this method has been used in this specific area of research. Table II presents the research questions that guided this review, and a description of the questions.
TABLE II: Research questions for the SLR process
No.
Research question
Description
RQ1
What localization and mapping approaches are used for the navigation of the visually impaired?
The target is to identify different localization and mapping techniques adopted for the development of assistive devices for visually impaired navigation.
RQ2
What are the advantages and limitations of SLAM techniques for BVI navigation?
The objective is to summarize the advantages and constraints of SLAM-based approaches for visually impaired navigation.
RQ3
What challenging situations have been addressed?
The purpose of this question is to know which of the challenging conditions (e.g. crowded environment, changing view point, challenging light conditions, dynamic objects, etc.) relevant to navigation systems have been considered.
RQ4
How does the proposed solution improve navigation using SLAM for individuals with impaired vision?
The research question seeks to understand how SLAM techniques can enhance mobility and navigation in individuals with visual impairments.
II-B2 Search strategy
A key step in performing SLR is to design an effective search strategy. This strategy should be executed with reasonable effort to retrieve relevant studies from digital libraries [54]. The exhaustive search process in systematic reviews is a critical factor distinguishing them from traditional literature reviews [54], leading to a wider and more precise understanding of the topic under review. To design the search string presented in Figure 2 we first extracted keywords from the PICOC elements, including population, intervention, and outcomes. We then determined synonyms for each keyword to broaden the search string. The list of keywords and their synonyms related to each PICOC element is listed in Table III.
Figure 2: Search string
TABLE III: Keywords used to design search string
Keywords
Synonyms
PICOC element
Visually impaired navigation
Blind navigation, Navigation assistance for the visually impaired, Navigational guidance for individuals with visual impairments, Navigational support for the visually disabled, Orientation and mobility for the visually impaired, Sight-impaired navigation, Visual impairment navigation aid, low vision navigation, partially sighted navigation
Population
SLAM
Real-time mapping and positioning, Simultaneous Localization and Mapping, Simultaneous mapping and position tracking, mapping and localization
Intervention
Accurate, efficient, reliable assistive technology
high-performance, precise, effective, trustworthy assistive technology
Outcome
The first part of the search string (i.e. ’visual* impair*’ OR ’blind’ OR ’visually disabled’ OR ’sight impaired’) is relevant to the population element of the PICOC framework. The addition of an asterisk to the terms ’visual’ and ’impair’ allows us to include various expressions, including ’visually impaired’ and ’visual impairment.’ The following segment of the query, consisting of ’navigation*’ OR ’mobility’ OR ’wayfinding,’ also places emphasis on the population aspect within the context of the systematic review. The inclusion of an asterisk in ’navigation*’ ensures comprehensive coverage, accounting for variations such as ’navigational.’ Regarding the Intervention component defined in the PICOC framework, we employed the term ’SLAM’ in conjunction with synonyms identified in the literature from diverse domains where SLAM is applied, such as robotics, autonomous driving cars, and underwater SLAM. The last segment of the search string is connected to the outcome element of the PICOC. Adding keywords such as ”localization” alone or the specific names of SLAM techniques did not increase the number of related papers. The search strings were employed on ten large citation databases, as listed in Table IV, to carry out an exhaustive search. We modified the base search string according to the Search Tip in each library to satisfy specific requirements.
TABLE IV: Databases selected for the search procedure
Digital source
Web address
# of papers
Last access date
ACM Digital Library
http://portal.acm.org
4
22 Jul 2023
Google Scholar
https://scholar.google.com
0
23 Jul 2023
IEEE Xplore
http://ieeexplore.ieee.org
11
22 Jul 2023
MDPI
https://www.mdpi.com
2
23 Jul 2023
PubMed
https://www.ncbi.nlm.nih.gov/pubmed
3
22 Jul 2023
Science Direct
http://www.sciencedirect.com
2
22 Jul 2023
Scopus
http://www.scopus.com
22
23 Jul 2023
Springer Link
http://link.springer.com
2
22 Jul 2023
Taley & Francis
https://www.tandfonline.com
0
22 Jul 2023
Wiley Online Library
https://onlinelibrary.wiley.com
1
22 Jul 2023
We utilized the Advanced Search feature in digital libraries to gain more control over our search parameters. The title, abstract, and keyword fields were selected to retrieve the search results. Searching on Google Scholar is somewhat different from searching for other digital libraries. Unlike other platforms, Google Scholar does not suggest various filters, requiring the manual incorporation of filters into the search string. Additionally, to identify English-written studies, we adjusted the language preference settings within our Google Scholar account to filter the search results in English. Upon receiving the ”Use fewer boolean connectors (max 8 per field)” message on Sciencedirect, we divided search string into separate queries to keep the number of boolean connectors within the limit. We then executed these queries separately, and merged the results to obtain a set of relevant studies.
II-B3 Selection criteria
Table V presents the selection criteria used to identify the eligible studies during the selection process. The Availability criterion included studies accessible in full text from digital databases. In addition, the Language criterion ensured the inclusion of publications written only in English. Furthermore, the Publication Period criterion restricted the inclusion of studies published between January 2017 and July 2023. This timeframe allowed us to prioritize the most current and state-of-the-art approaches. The Type of Source criterion included conference and journal papers, which were considered peer-reviewed and academically recognized sources. Books, dissertations, newsletters, speeches, technical reports, and white papers were excluded. Finally, the Relevance criterion played an important role in the exclusion process; therefore, publications outside the scope of our study were excluded based on a review of their titles and abstracts.
TABLE V: Selection criteria
Criteria
Inclusion
Exclusion
Availability
Available in full text
Not accessible in specific databases
Language
English
Not written in English
Publication period
From 2017 to July 2023
Prior 2017
Type of source
Conferences and journals papers
Books, dissertation, newsletters, speeches, technical reports, white papers
Relevance
Papers relevant to at least two research questions
Outside the scope of our research
II-C Conducting the review
As shown in Figure 1, the review process began after the review protocol was finalized. The conducting phase is a multi-stage process that includes research identification, study selection, data extraction, and data synthesis. In the research identification step, digital libraries were searched using adapted search strings that were specific to each library. This search aimed to collect a pool of potentially relevant primary studies. The next step involved the selection of studies for which the relevance of each study to the review was evaluated. The steps involved in this process are illustrated in Figure. 3. During the data extraction phase, the data required from the studies were collected and analyzed. We employed the data extraction form established during the development phase of the review protocol to ensure accurate extraction of information that addresses our research questions.
II-C1 Identification
During the initial phase of our review, we conducted searches across the digital libraries using custom-formulated queries for each library. The initial searches yielded a large number of papers irrelevant to our research questions, resulting in 6,809 records. As explained in Section II-B2, we employed a single query in all libraries, except ScienceDirect and Google Scholar, to retrieve relevant records. The results obtained from these searches were exported in the BibTex format, a process facilitated by the export citation features available in the libraries. The BibTex data were then imported into the Parsifal framework for subsequent stages of our review. Springer and Google Scholar do not provide direct options for exporting data in BibTex format. To address this issue, we used Zotero and its browser plugin, Zotero Connector, to streamline the process. With these tools, we added paper information from webpage views to Zotero and subsequently retrieved BibTex data.
For Springer Link, which provides only CSV files with search results, we opened the CSV in Excel and extracted the DOIs. These DOIs were then pasted into Zotero’s ”Add item(s) by identifier” feature. After importing the DOIs into Zotero, we selected the appropriate folder containing the imported papers and exported the collection to the BibTex format using a simple right-click. As Scholar Google does not provide easy export of a large number of records, we adopted a similar approach: creating a library, saving search results to that library, and exporting paper data in BibTex format from that library. This process ensured that we obtained the necessary data for the subsequent stages of our systematic review.
II-C2 Study selection
After conducting searches in the digital libraries, we applied our selection criteria, as defined in our review protocol, to filter out irrelevant studies. Initially, records published before 2017 were excluded. Further exclusions involved filtering out publications that were not written in English or had not been published in peer-reviewed venues. Following these steps, of the initial 6809 records found in the initial search, 5431 were excluded.
We imported the study data into the Parsifal platform in BibTeX format, as explained in Section II-C1, which helped remove duplicate studies. A total of 116 duplicate papers were excluded. We then reviewed the titles and abstracts of the remaining studies, excluding those irrelevant to our research topic. In this step, 779 studies were excluded.
In the next step, we performed a fast reading of the full text of the remaining papers, excluding 265 studies that were outside the scope of our research. We then evaluated the quality of the studies based on the quality assessment criteria defined in the SLR protocol. Five studies were removed during the assessment of study quality. Table VI lists the quality assessment criteria for our SLR.
TABLE VI: Quality assessment criteria and weights
Criteria
Weight
Is there an adequate description of the context in which the research was carried out?
0.0, 0.5, 1.0
Does the methodology take into account both localization and mapping issues?
0.0, 0.1, 2.0
Is the solution proposed well presented?
0.0, 0.5, 1.0
Is there a clear statement of findings?
0.0, 0.5, 1.0
Is the research design appropriate to address the aims of the research?
0.0, 0.5, 1.0
Does the study add value to the research community?
0.0, 0.5, 1.0
We carefully read 213 full-text articles to address the research questions. As 166 articles were not relevant to at least two of our research questions, they were removed, leaving 47 articles for the final stage.
To objectively assess the performance of our search strategy, we employed the quasi-gold standard (QGS) technique, as described by [54]. Using this method, a set of articles related to the research topic is manually selected. Digital libraries are then searched based on the research strategy to identify related studies. Finally, the retrieved articles are compared with QGS, and the sensitivity of the search strategy is calculated using the following formula:
[…]
In our SLR, with 30 manually selected relevant studies and 48 studies retrieved using the SLR search strategy, of which 26 were among the manually selected studies, the resulting sensitivity was approximately 86.67%.
To provide a broader range of relevant studies, we included papers on the forward snowballing process [55]. This process involves identifying and accessing references in a paper and reviewing cited papers. We used ”Cited by” feature of Google Scholar to identify these additional papers. In this stage, 695 articles were identified. After removing duplicates and applying selection criteria similar to those used for the articles obtained from digital libraries, we added seven more articles to the final collection. Consequently, 54 articles were included in this review. Figure 3 shows a diagram of the study-selection process.
Figure 3: Studies selection process
It is important to note that the last search conducted in digital libraries was on July 23, 2023, and that for forward snowballing was on August 12, 2023. These dates should be considered as the starting points for future reviews. The publications included in our review are listed in Tables VIIIX and categorized based on their publication venues. These tables provide an overview of the literature, including the paper title, author names, publication year, location, and source through which the studies were discovered. Among the 54 studies included in our SLR, 27 were sourced from journals, as presented in Table VII. The remaining 27 studies were presented at conferences, as shown in Tables VIII and IX.
II-C3 Data extraction
Data extraction is a critical phase in the systematic literature review process in which relevant data from selected studies are systematically collected. To achieve this objective, we employed the data-extraction form defined in the SLR protocol. This form consists of various fields designed to retrieve answers to our research questions from each of the included articles. Within the scope of this SLR, we defined the following essential elements, each contributing to a comprehensive understanding of the reviewed literature:
  • Short summary of the paper: A concise overview of the main points and findings of the study
  • Localization and mapping technique: Identification of specific techniques applied for localization and mapping.
  • Advantages of the presented method: The strengths associated with the localization method presented in each paper for visually impaired navigation.
  • Limitations of the presented method: Identification of weaknesses or constraints associated with the localization technique.
  • Types of obstacles addressed: Categorization of obstacles, static and dynamic, as a challenge during navigation.
  • Challenging conditions: Explanation of other challenging scenarios that the methods are designed to handle.
  • Types of sensors: Identification of sensors employed to receive data from the surroundings.
  • Improvement in navigation: Identification of how SLAM-based methods enhance navigation for individuals with impaired vision.
  • Possible future opportunities and directions: Exploration of potential future research areas and directions stemming from these findings.
  • The research questions addressed: Identification of specific SLR research questions addressed by each study.
  • Working area: Whether the method is intended for indoor, outdoor, or both indoor and outdoor environments.
TABLE VII: List of journal publications included in SLR
Ref. Title Authors Published Year Source
[56]
Sonification of navigation instructions for people with visual impairment
Dragan Ahmetovic and Federico Avanzini and Adriano Baratè and Cristian Bernareggi and Marco Ciardullo and Gabriele Galimberti and Luca A. Ludovico and Sergio Mascetti and Giorgio Presti
International Journal of Human-Computer Studies
2023
Science@Direct
[57]
Sensing and Navigation of Wearable Assistance Cognitive Systems for the Visually Impaired
Li, Guoxin and Xu, Jiaqi and Li, Zhijun and Chen, Chao and Kan, Zhen
IEEE Transactions on Cognitive and Developmental Systems
2023
IEEE Digital Library
[58]
Mixture reality-based assistive system for visually impaired people
Jucheng Song and Jixu Wang and Shuliang Zhu and Haidong Hu and Mingliang Zhai and Jiucheng Xie and Hao Gao
Displays
2023
Science@Direct
[59]
Research on Design and Motion Control of a Considerate Guide Mobile Robot for Visually Impaired People
Zhang, Bin and Okutsu, Mikiya and Ochiai, Rin and Tayama, Megumi and Lim, Hun-Ok
IEEE Access
2023
Scopus
[60]
UNav: An Infrastructure-Independent Vision-Based Navigation System for People with Blindness and Low Vision
Yang, Anbang and Beheshti, Mahya and Hudson, Todd E and Vedanthan, Rajesh and Riewpaiboon, Wachara and Mongkolwat, Pattanasak and Feng, Chen and Rizzo, John-Ross
Sensors
2022
PubMed
[61]
Multi-Floor Indoor Localization Based on Multi-Modal Sensors
Zhou, Guangbing and Xu, Shugong and Zhang, Shunqing and Wang, Yu and Xiang, Chenlu
Sensors
2022
Scopus
[62]
Knowledge driven indoor object-goal navigation aid for visually impaired people
Hou, Xuan and Zhao, Huailin and Wang, Chenxu and Liu, Huaping
Cognitive Computation and Systems
2022
Wiley Online Library
[63]
Indoor-Guided Navigation for People Who Are Blind: Crowdsourcing for Route Mapping and Assistance
Plikynas, Darius and Indriulionis, Audrius and Laukaitis, Algirdas and Sakalauskas, Leonidas
Applied Sciences (Switzerland)
2022
Scopus
[64]
A Multi-Sensory Guidance System for the Visually Impaired Using YOLO and ORB-SLAM
Xie, Zaipeng and Li, Zhaobin and Zhang, Yida and Zhang, Jianan and Liu, Fangming and Chen, Wei
Information
2022
MDPI
[65]
Egocentric Human Trajectory Forecasting with a Wearable Camera and Multi-Modal Fusion
Qiu, Jianing and Chen, Lipeng and Gu, Xiao and Lo, Frank P.-W. and Tsai, Ya-Yen and Sun, Jiankai and Liu, Jiaqi and Lo, Benny
IEEE Robotics and Automation Letters
2022
Scopus
[66]
A wearable navigation device for visually impaired people based on the real-time semantic visual slam system
Chen, Zhuo and Liu, Xiaoming and Kojima, Masaru and Huang, Qiang and Arai, Tatsuo
Sensors
2021
PubMed
[67]
Multimodal sensing and intuitive steering assistance improve navigation and mobility for people with impaired vision
Slade, Patrick and Tambe, Arjun and Kochenderfer, Mykel J.
Science Robotics
2021
Scopus
[68]
Assistive Navigation Using Deep Reinforcement Learning Guiding Robot With UWB/Voice Beacons and Semantic Feedbacks for Blind and Visually Impaired People
Lu, Chen-Lung and Liu, Zi-Yan and Huang, Jui-Te and Huang, Ching-I and Wang, Bo-Hui and Chen, Yi and Wu, Nien-Hsin and Wang, Hsueh-Cheng and Giarré, Laura and Kuo, Pei-Yi
Frontiers in Robotics and AI
2021
Scopus
[69]
Indoor Wearable Navigation System Using 2D SLAM Based on RGB-D Camera for Visually Impaired People
Hakim, Heba and Fadhil, Ali
Advances in Intelligent Systems and Computing
2021
Scopus
[70]
An RGB-D Camera Based Visual Positioning System for Assistive Navigation by a Robotic Navigation Aid
Zhang, He and Jin, Lingqiu and Ye, Cang
IEEE/CAA Journal of Automatica Sinica
2021
IEEE Digital Library
[71]
Hierarchical visual localization for visually impaired people using multimodal images
Cheng, Ruiqi and Hu, Weijian and Chen, Hao and Fang, Yicheng and Wang, Kaiwei and Xu, Zhijie and Bai, Jian
Expert Systems with Applications
2021
Scopus
[72]
Indoor Topological Localization Based on a Novel Deep Learning Technique
Liu, Qiang and Li, Ruihao and Hu, Huosheng and Gu, Dongbing
Cognitive Computation
2020
Scopus
[73]
Combining Obstacle Avoidance and Visual Simultaneous Localization and Mapping for Indoor Navigation
Jin, SongGuo and Ahmed, Minhaz Uddin and Kim, Jin Woo and Kim, Yeong Hyeon and Rhee, Phill Kyu
Symmetry
2020
MDPI
[74]
Wearable travel aid for environment perception and navigation of visually impaired people
Bai, Jinqiang and Liu, Zhaoxiang and Lin, Yimin and Li, Ye and Lian, Shiguo and Liu, Dijun
Electronics (Switzerland)
2019
Scopus
[75]
An ARCore based user centric assistive navigation system for visually impaired people
Zhang, Xiaochen and Yao, Xiaoyu and Zhu, Yi and Hu, Fei
Applied Sciences (Switzerland)
2019
Scopus
[76]
Virtual-Blind-Road Following-Based Wearable Navigation Device for Blind People
Bai, Jinqiang and Lian, Shiguo and Liu, Zhaoxiang and Wang, Kai and Liu, Dijun
IEEE Transactions on Consumer Electronics
2018
IEEE Digital Library
[77]
An indoor wayfinding system based on geometric features aided graph SLAM for the visually impaired
Zhang, He and Ye, Cang
IEEE Transactions on Neural Systems and Rehabilitation Engineering
2017
PubMed
[78]
Plane-Aided Visual-Inertial Odometry for 6-DOF Pose Estimation of a Robotic Navigation Aid
Zhang, He and Ye, Cang
IEEE Access
2020
Scopus
[79]
SRAVIP: Smart Robot Assistant for Visually Impaired Persons
Albogamy, Fahad and Alotaibi, Turk and Alhawdan, Ghalib and Mohammed, Faisal
International Journal of Advanced Computer Science and Applications
2021
Forward Snowballing
[80]
A Lightweight Approach to Localization for Blind and Visually Impaired Travelers
Crabb, Ryan and Cheraghi, Seyed Ali and Coughlan, James M
Sensors
2023
Forward Snowballing
[82]
Wearable system to guide crosswalk navigation for people with visual impairment
Son, Hojun and Weiland, James
Frontiers in Electronics
2022
Forward Snowballing
[83]
Indoor Low Cost Assistive Device using 2D SLAM Based on LiDAR for Visually Impaired People
Hakim, Heba and Fadhil, Ali
Iraqi Journal for Electrical & Electronic Engineering
2019
Forward Snowballing
TABLE VIII: List of conference papers included in SLR (part 1)
Ref. Title Authors Published Year Source
[84]
Efficient Real-Time Localization in Prior Indoor Maps Using Semantic SLAM
Goswami, R. G. and Amith, P. V. and Hari, J. and Dhaygude, A. and Krishnamurthy, P. and Rizzo, J. and Tzes, A. and Khorrami, F.
9th Inter. Conf. on Automation, Robotics and Applications (ICARA)
2023
IEEE Digital Library
[85]
Detect and Approach: Close-Range Navigation Support for People with Blindness and Low Vision
Hao, Yu and Feng, Junchi and Rizzo, John-Ross and Wang, Yao and Fang, Yi
European Conf. on Computer Vision
2022
Springer Link
[87]
PathFinder: Designing a Map-Less Navigation System for Blind People in Unfamiliar Buildings
Kuribayashi, Masaki and Ishihara, Tatsuya and Sato, Daisuke and Vongkulbhisal, Jayakorn and Ram, Karnik and Kayukawa, Seita and Takagi, Hironobu and Morishima, Shigeo and Asakawa, Chieko
CHI Conf. on Human Factors in Computing Systems
2023
ACM Digital Library
[88]
A Novel Perceptive Robotic Cane with Haptic Navigation for Enabling Vision-Independent Participation in the Social Dynamics of Seat Choice
Agrawal, Shivendra and West, Mary Etta and Hayes, Bradley
IEEE Inter. Conf. on Intelligent Robots and Systems
2022
Scopus
[89]
A Multi-Sensory Blind Guidance System Based on YOLO and ORB-SLAM
Rui, Chufan and Liu, Yichen and Shen, Junru and Li, Zhaobin and Xie, Zaipeng
IEEE Inter. Conf. on Progress in Informatics and Computing (PIC)
2021
IEEE Digital Library
[90]
Indoor Navigation Assistance for Visually Impaired People via Dynamic SLAM and Panoptic Segmentation with an RGB-D Sensor
Ou, Wenyan and Zhang, Jiaming and Peng, Kunyu and Yang, Kailun and Jaworek, Gerhard and Müller, Karin and Stiefelhagen, Rainer
Inter. Conf. on Computers Helping People with Special Needs
2022
Scopus
[91]
A Wearable Robotic Device for Assistive Navigation and Object Manipulation
Jin, Lingqiu and Zhang, He and Ye, Cang
IEEE Inter. Conf. on Intelligent Robots and Systems
2021
Scopus
[92]
Multi-functional smart E-glasses for vision-based indoor navigation
Xu, Jiaqi and Xia, Haisheng and Liu, Yueyue and Li, Zhijun
Inter. Conf. on Advanced Robotics and Mechatronics (ICARM)
2021
Scopus
[93]
Personalized Navigation that Links Speaker’s Ambiguous Descriptions to Indoor Objects for Low Vision People
Lu, Jun-Li and Osone, Hiroyuki and Shitara, Akihisa and Iijima, Ryo and Ryskeldiev, Bektur and Sarcar, Sayan and Ochiai, Yoichi
Inter. Conf. on Human-Computer Interaction
2021
Springer Link
[95]
Guiding Blind Pedestrians in Public Spaces by Understanding Walking Behavior of Nearby Pedestrians
Kayukawa, Seita and Ishihara, Tatsuya and Takagi, Hironobu and Morishima, Shigeo and Asakawa, Chieko
Proc. ACM Interact. Mob. Wearable Ubiquitous Technol.
2020
ACM Digital Library
[96]
A Navigation Aid for Blind People Based on Visual Simultaneous Localization and Mapping
Chen, Cing-Han and Wang, Chien-Chun and Lin, Sian-Fong
IEEE Inter. Conf. on Consumer Electronics
2020
IEEE Digital Library
[97]
Can we unify perception and localization in assisted navigation? an indoor semantic visual positioning system for visually impaired people
Chen, Haoye and Zhang, Yingzhi and Yang, Kailun and Martinez, Manuel and Müller, Karin and Stiefelhagen, Rainer
Computers Helping People with Special Needs: 17th Inter. Conf., ICCHP
2020
Scopus
[98]
Indoor Localization for Visually Impaired Travelers Using Computer Vision on a Smartphone
Fusco, Giovanni and Coughlan, James M.
17th Inter. web for all Conf.
2020
ACM Digital Library
[99]
Human-Robot Interaction for Assisted Wayfinding of a Robotic Navigation Aid for the Blind
Zhang, He and Ye, Cang
Inter. Conf. on Human System Interaction, HSI
2019
Scopus
[100]
A Multi-Sensor Fusion System for Improving Indoor Mobility of the Visually Impaired
Zhao, Yu and Huang, Ran and Hu, Biao
Chinese Automation Congress (CAC)
2019
IEEE Digital Library
[101]
Navigation Agents for the Visually Impaired: A Sidewalk Simulator and Experiments
Weiss, Martin and Chamorro, Simon and Girgis, Roger and Luck, Margaux and Kahou, Samira E. and Cohen, Joseph P. and Nowrouzezahrai, Derek and Precup, Doina and Golemo, Florian and Pal, Chris
Proceedings of Machine Learning Research
2019
Scopus
[102]
Real-time localization and navigation in an indoor environment using monocular camera for visually impaired
Ramesh, Kruthika and Nagananda, S. N. and Ramasangu, Hariharan and Deshpande, Rohini
5th Inter. Conf. on Industrial Engineering and Applications (ICIEA)
2018
IEEE Digital Library
TABLE IX: List of conference papers included in SLR (part 2)
Ref. Title Authors Published Year Source
[103]
Indoor Navigation using Text Extraction
Eden, Jake and Kawchak, Thomas and Narayanan, Vijaykrishnan
IEEE Inter. Workshop on Signal Processing Systems (SiPS)
2018
IEEE Digital Library
[104]
Autonomous Scooter Navigation for People with Mobility Challenges
Mulky, Rajath Swaroop and Koganti, Supradeep and Shahi, Sneha and Liu, Kaikai
IEEE Inter. Conf. on Cognitive Computing (ICCC)
2018
IEEE Digital Library
[105]
Localizing people in crosswalks using visual odometry: Preliminary results
Lalonde, Marc and St-Charles, Pierre-Luc and Loupias, Délia and Chapdelaine, Claude and Foucher, Samuel
Inter. Conf. on Pattern Recognition Applications and Methods (ICPRAM)
2018
Scopus
[106]
Plane-aided visual-inertial odometry for pose estimation of a 3D camera based indoor blind navigation
Zhang, He and Ye, Cang
British Machine Vision Conf. (BMVC)
2017
Scopus
[107]
CCNY Smart Cane
Chen, Qingtian and Khan, Muhammad and Tsangouri, Christina and Yang, Christopher and Li, Bing and Xiao, Jizhong and Zhu, Zhigang
IEEE 7th Annual Inter. Conf. on CYBER Technology in Automation, Control, and Intelligent Systems (CYBER)
2017
IEEE Digital Library
[108]
A Cloud and Vision-Based Navigation System Used for Blind People
Bai, Jinqiang and Liu, Dijun and Su, Guobin and Fu, Zhongliang
Inter. Conf. on artificial intelligence, automation and control technologies
2017
ACM Digital Library
[109]
Indoor positioning and obstacle detection for visually impaired navigation system based on LSD-SLAM
Endo, Yuki and Sato, Kei and Yamashita, Akihiro and Matsubayashi, Katsushi
Inter. Conf. on Biometrics and Kansei Engineering (ICBAKE)
2017
Scopus
[110]
SeeWay: Vision-Language Assistive Navigation for the Visually Impaired
Yang, Zongming and Yang, Liang and Kong, Liren and Wei, Ailin and Leaman, Jesse and Brooks, Johnell and Li, Bing
IEEE Inter. Conf. on Systems, Man, and Cybernetics (SMC)
2022
Forward Snowballing
[111]
The Methods of Visually Impaired Navigating and Obstacle Avoidance
Shahani, Siddharth and Gupta, Nitin
Inter. Conf. on Applied Intelligence and Sustainable Computing (ICAISC)
2023
Forward Snowballing
[112]
The Design of Person Carrier Robot using SLAM and Robust Salient Detection
Yun, Youngjae and Gwon, Taeyang and Kim, Donghan
18th Inter. Conf. on Control, Automation and Systems (ICCAS)
2018
Forward Snowballing
III Result
In this section, the findings of SLR are presented. Figure 4 shows the number of studies included in this review, which focused on BVI navigation using SLAM techniques. As shown in the figure, although only papers published in the first half of 2023 are included in this review, they constitute a substantial portion of the total. The growth in the number of studies in this domain suggests an advancement in SLAM techniques and an increase in their usage for developing navigation technologies for visually impaired individuals. This section is divided into four parts to answer the research questions. It discusses the types of SLAM techniques used to develop assistive technologies for visually impaired navigation, delves into the advantages and limitations of these techniques, highlights the challenging scenarios addressed, and presents the attributes of the SLAM technology that contribute to the enhancement of visually impaired navigation.
Figure 4: Publications included in this review on SLAM-based BVI navigation, by year
III-A RQ1. What localization and mapping approaches are used for the navigation of the visually impaired?
Since there are many different types of SLAMs designed with different types of sensors, for different applications, and for different scenarios, in this section, we focus on reviewing the types of SLAM that have been used for navigation of visually impaired people. It is a key technology in robotics and computer vision and has the potential to assist visually impaired individuals with navigation. This can help visually impaired individuals by providing real-time location information, maps, and spatial awareness. Among the 54 papers surveyed, three strategies were common, as shown in Table X. In this table, we use the exact terms mentioned in the literature for the localization and mapping techniques.
The majority of studies have leveraged established SLAM techniques, such as ORB-SLAM, while some studies have developed new solutions tailored to their needs. For example, [73] proposed visual simultaneous localization and mapping for the moving-person tracking (VSLAMMPT) method. The proposed method was designed to assist people with disabilities, particularly visually impaired individuals, in navigating indoor environments. Additionally, various studies have used the SLAM components of existing frameworks, such as ARCore and ZED camera SLAM.
It is worth noting that several studies have employed VIO and SLAM as the core components in their proposed systems, whereas others have employed them to enhance the robustness of localization [71], for comparison with alternative localization approaches [68], or to develop new localization methods [80]. For example, [80] presented a novel localization approach specifically designed for individuals with visual impairment. This method combines visual landmark identification, VIO, and spatial constraints derived from a two-dimensional (2D) floor plan.
SLAM methodologies can be categorized into feature-based, direct, and optical-flow techniques. Feature-based methods extract and describe feature points in an image, which are then matched across different images for tracking and mapping. On the other hand, direct methods directly calculate the luminosity changes of pixel blocks. Optical flow methods utilize the optical flow changes in feature points, pixel gradient points, or the entire image to track and map the environment [66].
SLAM algorithms are further categorized into optimization- and filtering-based methods, each with distinct approaches to map creation and agent localization. Optimization-based, often referred to as Graph SLAM, treats the problem as a large optimization task, where the goal is to find the set of poses and landmark positions that best explain the observed sensor measurements. This is typically achieved by constructing a graph in which nodes represent agent poses or landmarks and edges represent constraints or observations between them. The solution is determined by minimizing the global cost function, which represents the error between the predicted and actual measurements, using nonlinear optimization techniques. On the other hand, filtering-based SLAM uses recursive Bayesian filters, such as the Extended Kalman Filter (EKF) or Particle Filter, to incrementally update the map and the agent’s position as new sensor data arrive.
Systems such as ORB-SLAM and RTAB-Map are feature- and optimization-based, employing features and graph optimization for mapping and localization. Conversely, LSD-SLAM and DSO are examples of direct and optimization-based SLAM. Some systems, such as Semantic SLAM, may adopt either approach, depending on their implementation. It is important to note that the method and type of SLAMs are not directly mentioned in all papers, so the information provided here is a general categorization based on common practices within each category.
The reviewed studies demonstrate the versatility of SLAM in various navigation scenarios. The specific implementation and aspects addressed in each study varied depending on the application. SLAM can be employed in environments that lack a map and can dynamically create it while navigating. This involves simultaneous map creation and localization within an environment. Alternatively, SLAM can be employed to generate maps for subsequent navigation. In this case, SLAM first builds a map of the environment and then the map is utilized during navigation. Studies have also used SLAM odometry for navigational tracking. Odometry provides a continuous estimate of the position and orientation of an agent based on the sensor readings.
TABLE X: Localization and mapping approaches used for visually impaired navigation, with ”NA” denoting data not available.
Widely applicable techniques
Technique Type Method Sensor type Reference(s)
Cartographer
Scan matching
Optimization-based
LiDAR
[87, 59, 61, 95]
Hector SLAM
Scan matching
Optimization-based
LiDAR
[69, 112, 83]
LiDAR SLAM
Scan matching
Rao-Blackwellized particle filter
LiDAR
[79, 68, 67]
Dynamic-SLAM
Feature-based
Optimization-based
Visual
[90]
LSD-SLAM
Direct
Optimization-based
Visual
[109, 111]
OpenVSLAM
Feature-based
Optimization-based
Visual
[60, 93, 97]
ORB-SLAM
Feature-based
Optimization-based
Visual
[64, 89]
ORB-SLAM2
Feature-based
Optimization-based
Visual
[57, 62, 92, 101, 96, 76, 82]
ORB-SLAM3
Feature-based
Optimization-based
Visual
[63, 65]
Pose-graph SLAM
Feature-based
Optimization-based
Visual
[77]
RTAB-Map
Feature-based
Optimization-based
Visual
[104]
Semantic SLAM
NA
NA
Visual
[66, 84, 100]
Visual SLAM
Feature-based
Optimization-based
Visual
[85, 74, 102, 108, 58, 71, 72, 107]
DSO
Direct
Optimization-based
Visual
[105]
VIO (Visual Inertial Odometry)
Feature-based
Optimization-based
Visual
[56, 99, 98]
Customized solution
DVIO
Feature-based
Optimization-based
Visual
[70]
PAVIO
Feature-based
Optimization-based
Visual
[106, 78]
RGBD-VIO
Feature-based
Optimization-based
Visual
[91]
VSLAMMPT
Feature-based
Optimization-based
Visual
[73]
Spatial tracking frameworks
ARCore
Likely feature-based
Likely optimization-based
Visual
[75]
ZED camera’s SLAM
Likely feature-based
Not explicitly stated
Visual
[103]
ARKit
Likely feature-based
Not explicitly stated
Visual
[80, 110]
Intel RealSense SLAM
Likely feature-based
Not explicitly stated
Visual
[88]
Our analysis, underscored by the classifications in Table X, indicates a strong preference for feature-based and optimization-based SLAM approaches for visually impaired navigation. This preference is likely due to the robustness and efficiency of these methods in processing visual data, which is key for real-time assistive navigation. The common use of visual sensors in SLAM techniques can be attributed to advances in computer vision and image processing, which enhance navigation capabilities by providing rich environmental information. This makes visual-sensor-based SLAM techniques the most used in the implementation of assistive technologies for the blind, offering a cost-effective, versatile, and accurate solution for real-time navigation and spatial awareness.
Figure 5 provides insight into the use of various SLAM techniques for visually impaired navigation from 2017 to the date when SLR was conducted (July 2023). This figure illustrates that visual SLAM has consistently been used across all years. Although many other techniques also operate based on visual data, we mentioned each of these techniques, as indicated in the referenced studies.
The utilization of semantic SLAM and Cartographer SLAM signifies a recent trend towards leveraging advanced spatial understanding and mapping capabilities for visually impaired navigation. Semantic SLAM incorporates higher-level scene interpretation and enhances users’ contextual awareness. On the other hand, Cartographer SLAM provides SLAM in 2D and 3D across various platforms and sensor configurations, offering innovative solutions to tackle the diverse challenges associated with BVI navigation.
ORB-SLAM algorithms, including ORB-SLAM (published in 2015), ORB-SLAM2 (published in 2017), and ORB-SLAM3 (published in 2021), have gained popularity because of their robustness and performance. This can be attributed to its efficient feature extraction and matching techniques, making it well suited for real-time navigation applications.
Customized techniques have been developed to meet specific needs. This trend indicates that researchers have adjusted the SLAM techniques to better match the specific requirements of their intended applications. This suggests closer integration of SLAM with domain-specific needs.
Figure 5: Evolution and adoption of SLAM techniques in BVI navigation systems over time.
III-B RQ2. What are the advantages and limitations of SLAM techniques for BVI navigation?
III-B1 Advantages
Unlike many localization approaches, such as RFID- or GPS-based methods, which require infrastructure setup, SLAM does not depend on pre-existing infrastructure. It operates autonomously by creating maps and understanding the surroundings in real-time. SLAM relies on data captured by sensors already present on many mobile devices, such as smartphones, and offers a cost-effective solution for accurate localization and mapping without the need for additional hardware or subscription services.
One of the most important advantages of SLAM is its potential for real-time positioning, which determines the agent’s current location and orientation. Systems leveraging ORB-SLAM, for instance, excel in pose estimation by integrating various data types, including visual, inertial, and depth information, thus enhancing accuracy beyond conventional methods [64]. This feature is pivotal not only for effective navigation but also for obstacle detection and avoidance, ensuring the safety and confidence of visually impaired users as they navigate through immediate environments [64, 89].
SLAM’s ability to reuse and update maps incrementally allows for a high degree of environmental adaptation. Its capacity to relocalize within prebuilt maps or expand them as necessary ensures that users can rely on updated information for navigation [75, 76]. This adaptability is further enhanced by the capability of the system to handle dynamic environments, making it invaluable for visually impaired users who require real-time path adjustments in response to moving obstacles [73, 90].
Detailed environmental mapping facilitated by SLAM, ranging from two-dimensional layouts to complex 3D geometric and semantic maps, provides comprehensive spatial understanding [59, 67, 72]. Environmental awareness is critical for path planning and collision avoidance. Furthermore, the integration of semantic mapping enriches spatial understanding by adding contextually rich information to maps, thereby facilitating more informed decision-making and interaction with the environment [66, 100].
The integration of different types of sensors and technologies with SLAM significantly expands the scope of its application. By integrating techniques such as object detection algorithms or combining RGB-D and IMU sensor data, SLAM systems achieve a multilayered perception of the environment [64, 84]. Sensor fusion enhances a system’s ability to detect and classify objects, accurately navigate, and handle dynamic elements within the environment, thereby offering a more holistic assistive solution [83, 89].
The cost-effectiveness of SLAM-based solutions attributed to their reliance on widely available low-cost sensors makes this technology particularly interesting. Systems employing monocular cameras or wearable RGB-D cameras exemplify how SLAM can be implemented in a cost-effective manner without compromising functionality, thus making advanced navigational aids accessible to more users [69, 85]. The advantages of SLAM, derived from the literature, are listed in Table XI.
TABLE XI: Key features of SLAM that benefit visually impaired navigation.
Advantages Description Ref.
Accurate localization
SLAM provides precise positioning within an environment, essential for effective navigation and obstacle avoidance.
[56, 57, 58, 62, 64, 67, 68, 69, 70, 73, 74, 75, 76, 77, 78, 80, 82, 83, 85, 88, 89, 90, 91, 92, 93, 95, 97, 98, 99, 102, 105, 106, 107, 108, 109, 110, 112]
Environmental mapping
SLAM constructs maps of surroundings, enabling spatial awareness for navigation.
[57, 58, 59, 60, 61, 64, 67, 69, 70, 72, 73, 74, 75, 76, 77, 79, 83, 84, 87, 88, 89, 90, 92, 93, 96, 97, 100, 104, 107, 108, 109, 111, 112]
Map reuse
Previously created maps can be reused to enhance the efficiency and reduce the need for constant remapping.
[57, 75, 76, 89, 93, 109]
Loop closing
SLAM corrects trajectory drifts by recognizing previously visited locations, thereby improving long-term accuracy.
[57, 63, 84]
Semantic mapping
SLAM integrates contextual information into maps, thereby enhancing the understanding of the environment and its elements.
[66, 84, 100]
Object localization
SLAM identifies and positions objects within the environment, facilitating interaction and navigation around the obstacles.
[72, 85, 88, 91]
Dense navigation maps
SLAM generates detailed maps that provide rich environmental data that are crucial for complex navigation tasks.
[64, 66, 89]
Incremental map updating
SLAM continuously updates maps with new information, ensuring that they remain accurate and current.
[64, 69, 89]
Integration with other technologies
SLAM can be combined with other technologies and algorithms to enhance functionality and performance.
[64, 68, 72, 83, 84, 85, 89, 92, 96, 98]
Integration of sensors
SLAM utilizes a variety of sensors to enrich the environmental perception and mapping accuracy.
[58, 79, 84, 91]
Cost-effectiveness and accessibility
SLAM’s reliance on commonly available sensors makes it an affordable solution for widespread use.
[66, 69, 76, 83, 85]
Ground-truth trajectory
SLAM delivers accurate path tracking, aiding the development of reliable navigation instructions.
[63, 65, 76]
Integration with 2D map
SLAM data can be integrated with 2D maps to enhance the navigation accuracy and functionality.
[63, 70, 75, 77, 80, 98]
Dynamic environment handling
SLAM can be adapted to changes within environment, maintaining reliable navigation in the presence of moving obstacles.
[73, 90, 109]
Re-localization
SLAM can quickly regain accurate positioning after temporary tracking loss, ensuring continuous and reliable navigation.
[57]
III-B2 Limitations
Although SLAM technologies show great potential for improving navigation aids for the visually impaired, they are not without their limitations. These limitations can significantly impact the effectiveness and reliability of SLAM-based assistive systems.
A notable challenge is the computational complexity and the associated demand for system resources. The implementation of advanced SLAM algorithms and the integration of deep learning frameworks for semantic understanding introduce significant computational overhead [57, 72]. This complexity can compromise the real-time performance, which is crucial for assistive navigation. The need for appropriate hardware to process high-resolution data further underscores this limitation, potentially restricting the deployment of SLAM-based systems [88, 89].
The effectiveness of SLAM is dependent on its environmental characteristics. Accurate mapping and localization depend on the presence of distinct geometric features. In environments lacking such features or dynamically changing settings, SLAM systems may struggle to maintain accurate localization, thereby leading to navigation errors [77, 91]. This limitation is particularly evident in feature-poor areas such as long corridors or spaces with uniform surfaces, where loss of localization can occur [62].
Another critical limitation is the dependency on initial data or pre-existing maps. Some SLAM systems require sighted individuals to pre-map the environment, which can limit the flexibility and immediate usability of unmapped or altered spaces [75, 76]. This reliance on prior mapping can be a significant hurdle for deploying SLAM-based navigation aids in diverse and changing environments [82, 96].
Drifting errors present a substantial challenge for maintaining the long-term accuracy of SLAM systems. Over time, small inaccuracies can accumulate, leading to significant deviations from the true trajectory, which can disorient users and compromise the navigation safety [56]. In addition, the ineffectiveness of some SLAM systems for generating dense navigation maps limits their utility in providing the detailed guidance required for visually impaired navigation, necessitating further algorithmic enhancements [64].
The performance of SLAM in dynamic environments, characterized by moving obstacles and changing conditions, remains a critical concern. Systems may fail to adapt quickly to such changes, leading to potential navigation errors and safety risks for visually impaired users [68].
Some SLAM applications require external calibration or setup, such as placement of calibration boards in specific environments. This requirement can limit the spontaneity and ease of use of SLAM-based navigation aids because it imposes additional constraints [105]. Table XII outlines the overall limitations of SLAM, derived from the publications under review.
TABLE XII: Limitations of SLAM for visually impaired navigation.
Limitations Description Ref.
Complexity and computational requirements
SLAM’s advanced algorithms can demand significant computational power, impacting real-time performance and efficiency.
[57, 72, 79, 88, 89, 91]
Dependency on environmental features
The accuracy of SLAM depends on the presence of distinct environmental features, which limits its effectiveness in feature-poor or dynamically changing environments.
[77, 91, 111]
Dependence on initial data/prior maps
Some SLAM systems require pre-mapped environments or initial data setups by sighted individuals, thereby reducing the flexibility in unmapped or altered spaces.
[75, 76, 82, 89, 96]
Loss of localization in feature-poor areas
SLAM may experience frequent localization losses in areas lacking sufficient feature points, such as blank corridors or plain walls, thereby compromising navigation reliability.
[62]
Dependency on external calibration
The need for external calibration in certain SLAM applications can limit their spontaneity and practicality in unprepared environments.
[105]
Drifting error
Accumulating drifting errors in SLAM can reduce the long-term accuracy, leading to potential navigation inaccuracies and user disorientation.
[56]
Lack of dense navigation maps
The inability of certain SLAM systems to generate detailed dense maps can restrict their effectiveness in providing a comprehensive navigation guidance.
[64]
Vulnerability in dynamic environments
SLAM systems may struggle to adapt to dynamic environments with moving obstacles, thereby posing navigation challenges and safety risks.
[68]
III-C RQ3. What challenging situations have been addressed?
Optimal pathfinding, perception of the surroundings, and obstacle avoidance are crucial for navigation. The precise localization of a visually impaired user within the environment is essential for these functions to operate effectively. Our surveyed papers addressed the localization and mapping problems using various techniques. Some of these studies explored other navigation-related challenges. We categorized these challenges into two groups: those relevant to environmental complexities, and those related to the sensors used to receive environmental data. Dynamic obstacles and crowded spaces constitute the challenges in the first group, whereas challenges related to changing lighting conditions and the rapid motion of users that results in motion blur fall into the second group. Table XIII lists studies that investigated these challenges through the integration of SLAM with other approaches. In the following section, we discuss the studies that address these challenges.
TABLE XIII: Challenges addressed in reviewed studies using SLAM techniques
Related to Challenges Reference(s)
Environment
Crowded places
[59, 65, 68, 95]
Dynamic object
[59, 90, 68, 73, 109, 111]
Sensor
Fast motion
[72]
Illumination
[61, 71, 72, 74, 81]
III-C1 Environmental complexities
Crowded scenarios Navigating crowded environments presents significant challenges for the visually impaired, leading to increased collision risks and difficulties in maintaining personal space. The absence of visual information makes it difficult to measure the distance, perceive crowd density, and locate landmarks or places of interest.
Navigation in crowded environments also poses challenges for assistive technology. For example, in assistive systems that operate based on SLAM, the presence of numerous dynamic elements, such as moving individuals and objects, introduces ambiguity into feature detection and tracking, leading to difficulties in accurately estimating the pose of the user and structure of the environment. The dynamic nature of crowds also hinders loop closure detection, disrupts map consistency, and contributes to drifts. Moreover, the lack of distinct visual landmarks in crowded scenes represents a reliable localization challenge that potentially reduces the robustness and accuracy of the overall SLAM system. Addressing these challenges requires the development of new approaches to address the complexity of such environments effectively. Several studies have examined this issue.
[59] presented a guide mobile robot engineered for the complexities of navigating different environments while considering dynamic objects and human presence. The robot could handle crowded environments with multiple dynamic objects. To accomplish this, the robot leveraged a spatial risk map, a tool that evaluates potential object-occupied spaces, to chart a path that effectively minimizes disruptions. This paper presents experiments in which a robot successfully guided a user through the passage of multiple objects and people.
In another study, [65] introduced an egocentric human trajectory forecasting model designed for navigation in crowded environments. By leveraging a wearable camera, the model captures detailed surroundings and the user’s past trajectory to predict their future. The trajectories obtained using ORB-SLAM3 were used to train the egocentric human trajectory forecasting model, which predicted the wearer’s future trajectory based on the contextual cues captured by the wearable camera and the wearer’s past trajectory.
Additionally, [68] addressed challenges in crowded environments using a combination of SLAM and Ultra-Wideband (UWB) positioning. However, the SLAM algorithm was found to be less effective in environments with dynamic obstacles such as pedestrians. The algorithm finds features of dynamic obstacles moving along with the robot as the assistive device, and thus it was misled that the robot did not move at all. However, using UWB positioning mitigates this issue.
In [95], the method addressed crowded environments by recognizing and predicting people’s behavior while anticipating the collision risk. The system advises users to adjust their walking speed (on-path mode) or to choose alternative routes (off-path mode). This involves comparing the 3D point cloud map to the real-time LiDAR and IMU sensor data. The system then predicts the future position and velocity of the user in order to avoid collisions.
Dynamic objects Several publications have not directly addressed the challenge of crowded environments, and have only focused on dealing with the presence of dynamic objects within the scene. In the system proposed by [90], dynamic objects can be identified, and average depth information can be provided to the user. When a dynamic object belongs to a predefined class such as a person, it can also be tracked between frames in the SLAM pipeline. The system is capable of identifying and tracking dynamic objects after ego-motion estimation to obtain average depth information. Subsequently, it can estimate the poses and speeds of these tracked dynamic objects and relay this information to the users through acoustic feedback. Depth information helps users maintain social distancing in public indoor environments such as shopping malls. To address the challenge of dynamic objects, [73] proposed a new method called visual simultaneous localization and mapping for moving person tracking (VSLAMMPT). This method was designed to handle dynamic environments in which objects constantly move. The system also uses the expected error reduction with active-semi-supervised learning (EER–ASSL)-based person detection to eliminate noisy samples in dynamic environments. This aids in accurate detection and avoidance of dynamic obstacles.
III-C2 Sensor-related
Changes in lighting condition Lighting changes pose a hurdle to visual SLAM systems. Illumination variations alter the visual features, interfere with accurate detection and matching across frames, impact pose estimation, and map-building robustness. SLAM relies on distinctive features for operation; however, lighting changes introduce ambiguities, noise, and errors, which affect accuracy. Overcoming this challenge requires robust algorithms for dynamic lighting to ensure stable and precise localization and mapping.
The method proposed in [71] tackles the challenge of illumination changes using a deep descriptor network called a Dual Desc, which is designed to be robust against various appearance variations including illuminance changes. The network uses multimodal images (RGB, Infrared, and Depth) to generate robust attentive global descriptors and local features. These descriptors were used to retrieve coarse candidates of query images, and 2D local features, along with a 3D sparse point cloud, were used for geometric verification to select the optimal results from the retrieved candidates. The authors mentioned that their dataset included images captured at different times of the day, resulting in illumination changes between the query and database images. Despite these changes, the proposed method achieved satisfactory localization results.
The authors of [72] evaluated the influence of lighting conditions on the performance of their novel localization method. The authors captured training images during the day and test images at night, and simulated changes in lighting conditions by switching some of the lights off in locations without windows. The results showed that changes in the lighting conditions had a minor impact on the proposed method. [61] mentioned that the proposed localization scheme was verified in a typical office building environment with dramatically changing lighting conditions throughout the day; however, it does not provide detailed results or discussion on how changing lighting conditions affect the performance of the system.
The method proposed in [81] addressed changes in illumination as a challenge using the COLD and IDOL datasets, which were recorded under different weather and illumination conditions (cloudy, night, and sunny) using different mobile platforms and camera setups. These datasets were used to evaluate the strength of the localization and recognition algorithms with respect to the variations caused by human activities and changes in illumination conditions. The study also mentioned the use of Histogram of Oriented Gradients (HOG) for feature extraction, which provides preferable invariant results for lighting or shadowing.
Motion blur Blurred images in visual SLAM can lead to inaccuracies in feature detection and matching, causing issues with pose estimation, map-building, loop closure, and visual odometry. These inaccuracies can also impact depth measurements and map quality. To address this issue, strategies such as using high-frame-rate sensors, incorporating IMUs, and employing motion deblurring techniques can be employed to improve the accuracy of localization and mapping in SLAM systems. Motion blur can be caused by fast or sudden movements of the user during navigation, which can affect localization performance. [72] studied this challenge and evaluated the robustness of the localization methods. They captured 2316 blurred images on the testing day. The results showed that the proposed method performed poorly in this experiment, indicating that fast motion or sudden changes in user movement can pose a challenge to the system. The reason for this poor performance was that the object detection scores did not exceed the threshold during the experiment.
III-D RQ4. How the proposed solution is expected to enhance mobility and navigation for visually impaired?
This section discusses how the approaches proposed by the studies included in our SLR have the potential to improve navigation for BVI people. These studies focused on diverse attributes, such as accurate pose estimation, semantic mapping, sensor fusion, and algorithmic innovations, to improve the quality of BVI navigation. Table XIV presents the categorization of attributes that contribute to enhancing the mobility and navigation of visually impaired individuals.
SLAM technology is primarily used to provide precise localization, which is critical for assistive navigation systems. Precise localization provides accurate information regarding a user’s position in the environment in which the user navigates. This accuracy enables the system to offer feedback on obstacles, pathways, and points of interest, thereby allowing BVI to navigate safely and confidently. Real-time assistance has also emerged as the key feature. Providing immediate feedback on the environment enables users to travel efficiently and safely, which leads to increased mobility and independence. Semantic mapping generates maps beyond geometric data. Such representations contain not only spatial information but also the semantic meanings of objects and features within the environment. This semantic understanding offers a deeper insight into the environment. This contextual awareness is particularly beneficial for enhancing navigation accuracy, as it enables navigation systems to make decisions based on semantic context, improving obstacle avoidance, path planning, and overall navigation efficiency.
By employing robotic systems such as small robots, smart canes, and sensor-equipped suitcases, some studies have provided guidance, obstacle avoidance capabilities, and increased spatial awareness, thereby effectively providing independent navigation for visually impaired individuals. Smartphone-based solutions harness the ubiquitous nature of smartphones that are equipped with cameras and sensors. These solutions offer navigation assistance by using widely available and familiar devices. Both indoor and outdoor navigation capabilities offer a seamless transition between environments, ensuring that users receive consistent support, regardless of the scene in which they navigate. Innovative localization and mapping algorithms enhance navigation efficiency and effectiveness through tailored modifications of existing SLAM frameworks or through the creation of novel solutions. Ultimately, these advancements have led to an improved overall experience for individuals with visual impairment. Although these studies focused on different features and attributes, they all aimed to enhance mobility, independence, and overall quality of life for BVI people.
TABLE XIV: Attributes of SLAM-based navigation systems that contribute to enhancing BVI navigation, along with referenced papers emphasizing each feature.
Features Description Reference(s)
Precise Localization
Using SLAM algorithms, these solutions accurately estimate the position and orientation of visually impaired users. Precise localization is essential for visually impaired navigation systems in order to ensure accurate real-time guidance, obstacle avoidance, and spatial awareness, ultimately enhancing independent and safe mobility.
[109, 97, 95, 93, 90, 82, 78, 77, 76, 75, 73, 70, 69, 64, 89, 67]
Real-time assistance
This feature ensures that BVI users receive immediate feedback about their environment. Therefore, they can be guided to navigate safely and efficiently, thereby enhancing their overall mobility and independence.
[85, 57, 92, 82, 77, 69, 66, 67, 88]
Semantic mapping
Semantic mapping involves creating detailed environmental representations that go beyond geometric data and enables visually impaired users to navigate with a deeper understanding of their surroundings.
[100, 97, 66, 72, 84]
Both indoor and outdoor navigation
Systems that serve both indoor and outdoor settings allow users to transition seamlessly between different environments while receiving consistent support.
[85, 108, 67, 74, 111]
Innovative algorithms
Innovative algorithms lead to advancements in navigation techniques. These approaches contribute to a more efficient and effective navigation, ultimately improving the overall experience of the visually impaired.
[77, 66, 106, 78, 91, 73]
Robotic navigation
These solutions employ robotic systems, such as robots, smart canes, and suitcases, which are equipped with sensors to assist BVI in navigating their environment.
[112, 107, 78, 70, 64, 89, 67, 59, 68, 87]
Smartphone based
Equipped with cameras and sensors, smartphones can be considered versatile navigation tools. These solutions offer assistance through devices that are widely available and familiar.
[98, 80, 110, 109, 108, 107, 102, 77, 75, 63, 74]
IV Future opportunities
An effective navigation system for BVI people needs to meet mobility metrics, such as decreasing navigation time, decreasing navigation distance, decreasing contact with the environment, and increasing walking speed. These systems must be highly accurate and efficient in complex situations, such as crowded places and changing light and weather conditions. At the same time, assistive aids should be comfortable, easy to use, unobtrusive, cost-effective, lightweight, and reduce cognitive load.
In this review, we examined publications that employed SLAM techniques in their navigation approaches. One of the distinct advantages of SLAM is its applicability in diverse locations without the need for pre-built maps or additional infrastructure such as Bluetooth beacons or RFID tags. However, there is still room for improvement in various aspects of these systems, including their ability to handle complex scenarios, provide accurate obstacle information, and seamlessly transition between indoor and outdoor environments. This section discusses the open problems and research directions identified during the SLR.
Challenge scenarios and real-world studies Navigating crowded environments is a significant challenge for the visually impaired. However, studies included in the SLR that address this issue are limited. Among the studies that considered this challenge, evaluations were conducted in controlled settings, rather than addressing the complexities of real-world scenarios. It is essential to note that the applicability of these solutions to real-world scenarios such as train stations is crucial. Therefore, there is a pressing need for future studies focusing on the complexities of navigation in high-traffic public places. Similarly, confronting challenging conditions, such as changes in illumination, low-light scenarios, high-speed dynamic objects, and complex backgrounds, can enhance the robustness and versatility of navigation systems. To overcome these challenges in the field of visually impaired navigation, we can draw inspiration from other domains and adapt the proposed solutions. For instance, [116] introduced a system that utilized an image enhancement technique for ORB points and LSD line feature recovery. This approach enables the system to perform effectively in a broader range of scenarios, providing extensive spatial information, even in low-light and hazy agricultural environments.
Long-term navigation The development of solutions that are effective over extended navigation periods is critical for achieving autonomous navigation. These solutions must ensure accurate mapping and localization even when the maps are updated over a longer navigation duration. To address this challenge, researchers can leverage the proposed solutions in robotics, such as that presented in [117], which introduced a novel long-term SLAM system with map prediction and dynamic removal, allowing wheelchair robots to maintain precise navigation capabilities over extended periods.
Deep learning integration The integration of deep learning with the SLAM algorithms, adopted for BVI navigation, deserves further attention. Deep learning offers a versatile approach for enhancing various aspects of SLAM, such as precise pose estimation under challenging conditions, relocalization, and loop-closure detection. Despite challenges, such as the need for large, accurately labeled datasets, the black-box nature of deep-learning models, and the computational intensity of SLAM systems, the association between deep learning and SLAM holds promise for advancing navigation solutions for the visually impaired, particularly in challenging scenarios.
Indoor and outdoor navigation integration The need for navigation systems capable of seamless transitions in both indoor and outdoor environments is crucial for enhancing the independence and mobility of BVI individuals. However, most of the reviewed papers in our SLR have primarily concentrated on indoor environments. Future research could address this challenge by developing solutions that offer unified and consistent navigation experience.
Obstacle detection Achieving detailed knowledge of obstacles and their characteristics is essential for BVI people. Although some papers included in the SLR addressed obstacle detection, the depth and accuracy of the obstacle information provided may still be limited. Future research opportunities lie in enhancing the SLAM algorithms to provide context-aware obstacle information. Researchers can draw inspiration from approaches adopted in robotics and autonomous drones such as real-time metric-semantic SLAM, as demonstrated in [118].
Semantic information integration Integrating semantic information into SLAM algorithms can significantly enhance the performance and robustness of navigation systems for BVI individuals. This information can be used to refine the mapping and localization processes and enhance the overall reliability of navigation in complex environments. For instance, semantic information can help reject outliers in loop closure detection, which is a crucial step in SLAM that involves identifying and matching previously visited locations.
Product development and collaboration Notably, all the reviewed approaches were prototypes in the early stages of research and are not yet practical products. This might be due to the absence of a unified community or group dedicated to solving the BVI navigation challenges. Much of the work in this domain is carried out by academic groups or small companies that often fail to produce feasible final products. This underscores a significant future opportunity to develop collaboration and to bridge the gap between research and practical implementation.
In conclusion, the future of SLAM for the visually impaired navigation is promising. Continued research efforts have the potential to develop SLAM algorithms tailored for BVI navigation, empowering visually impaired individuals with a safe and independent means of navigating their surroundings.
V Conclusion
This study presents a systematic literature review of recent studies on SLAM-based solutions for BVI navigation. Excluding papers published before 2017, this review focused on the latest advancements, innovations, and considerations, resulting in a more relevant and comprehensive understanding of the current state of research. Relevant data were extracted from 54 selected studies that adhered to the SLR selection criteria to address the research questions. By analyzing the selected papers based on their SLAM techniques, we observed that the majority of the studies utilized visual SLAM techniques, such as ORB-SLAM3, owing to their advantages for visual sensors.
Several studies have introduced novel strategies for addressing localization and mapping challenges tailored to the specific requirements of their research, whereas certain studies have employed existing spatial tracking frameworks to develop navigation solutions. We also investigated the advantages and limitations of the SLAM techniques, as highlighted in the studies under review. Notably, most studies have leveraged accurate localization features of SLAM.
We also investigated the challenging scenarios encountered by SLAM-based navigation systems, which have been addressed in the literature. Furthermore, we discuss how the proposed SLAM-based solutions are expected to improve the mobility and navigation of the visually impaired. Finally, we identified gaps, opportunities, and areas of interest that could be explored further in future research, such as addressing challenges in crowded environments, improving real-world applicability, integrating deep learning, and ensuring long-term navigation effectiveness in SLAM-based solutions for visually impaired navigations.
Given the widespread application of SLAM in robotic, autonomous drones, and auto-driving car navigation, these techniques can be adapted to ensure safe and independent BVI navigation. This is particularly important in dynamic and challenging environments, including those with varying lighting conditions where research opportunities remain abundant. The potential of integrating these techniques into the navigation of visually impaired individuals continues to be an open and promising avenue.
Acknowledgment
We would like to extend our sincere gratitude to Giovanni Cioffi, whose insightful feedback and thorough review significantly contributed to the refinement of this manuscript.